OpenStudy (anonymous):

PLEASE!!!!! Solve the equation on the interval [0, 2π). sec X/2 = cos X/2?

OpenStudy (anonymous):

sec(x) is the inverse of cos(x)

OpenStudy (anonymous):

so you can put 1/cos(x/2) in place of sec(x/2) in the equation

OpenStudy (solomonzelman):

\[\sec(x/2)=\cos(x/2)~~~->~~\frac{1}{\cos(x/2)}=\frac{\cos(x/2)}{1}~~->~~\cos(x/2)=1\]can you take it from here?

OpenStudy (anonymous):

IS IT (0)?

OpenStudy (anonymous):

No you need to solve the equation cos^2(x/2) = 1

OpenStudy (anonymous):

\[\pi/4,5\pi/4?\]

OpenStudy (anonymous):

Yes it is pi/4 and 3pi/4

OpenStudy (solomonzelman):

cos^2(x/2) = 1 is same as cos(x/2) = ±1 this plus minus I left out, because 1/-1=-1/1, just like 1/1=1/1

OpenStudy (anonymous):

SO IM RIHT OR NOT/

OpenStudy (anonymous):

Your answer is half right pi/4 is right but how did you get 5pi/4 did you mean 3pi/4 ?

OpenStudy (anonymous):

IT WASNT NONE OF MY CHOICES I GOT 3PI/4 BUT IT WASNT IN MY CHOICES